OpenStudy (ashlyhall8):

The hypotenuse of a right triangle is 17. The lengths of the two legs are whole numbers. What are they?

zepdrix (zepdrix):

Is there some fancy way to do this? I'm not really sure. I'm sure we can figure it out the long way though, \(\large\rm a^2+b^2=c^2\) \(\large\rm a^2+b^2=17^2\) \(\large\rm a^2+b^2=289\) And from there we can just test out some numbers.

OpenStudy (biffo):

@zepdrix -- and from there, if we're permitted to go so far as the answer, we look up (or, if possible, remember) our list of Pythagorean triplets (often, sadly, called triples) and see that the adjacent sides have lengths of 8 and 15 units. Hope that doesn't spoil the fun too much.